### v4

@@ 6 Conclusions
This paper has identified two leading algorithms for solving kernel ridge regression (KRR) problems with a moderate number of data points ($10^{4}\leq N\leq 10^{7}$). The proposed algorithms based on RPCholesky and KRILL preconditioning are relatively fast, robust, and reliable, outperforming previous approaches [16, 22, 23, 53, 44, 45, 30].
@@
Eigenvalue decay is the main requirement currently limiting the performance of RPCholesky preconditioning, and this limitation becomes especially apparent when the regularization $\mu$ is very small (see the right panel of fig. 1). Future work is needed to discover effective preconditioners for KRR problems with slow eigenvalue decay. Existing methods based on hierarchical decompositions [16, 3, 11, 56] are a promising candidate for this purpose, but there are concerns these methods perform poorly in high dimensions [3].
@@
With appropriate parameter choices and neglecting rounding errors, KRILL is guaranteed to control the condition number regardless of the regularization and the kernel matrix eigenvalue decay. Despite this success, there remain opportunities to improve on KRILL. First, the existing analysis of sparse random sign embeddings [15] does not specify appropriate values of $\zeta$ and $d$ to use in practice. More analysis of sparse sign embeddings is needed to close the theory–practice gap; the recent paper [13] makes some progress in this direction. A second potential concern with KRILL is numerical stability; see the recent papers [32, 21]. As a practical remedy, we have addressed numerical stability by adding a small shift to the regularizer, but this approach has not been analyzed rigorously. Finally, it may be possible to develop more efficient algorithms for restricted KRR problems when the number of data centers is $k\gg N^{1/2}$. The ideal conjugate gradient algorithm would require $\mathcal{O}(kN)$ operations, but KRILL is more expensive due to the cost of inverting the $k\times k$ preconditioner. The block Lanczos preconditioner discussed in Section 3.2 may be interesting in this setting, though—unlike KRILL—it requires spectral decay to be effective.
@@ Acknowledgments
We would like to acknowledge helpful conversations with Misha Belkin, Tyler Chen, Riley Murray, Madeleine Udell, and Jorge Garza-Vargas.
@@ Appendix A Data sets
Data set Dimension $d$ Sample size $N$ Source

### v5

@@ 6 Conclusions
This paper has identified two algorithms for solving kernel ridge regression (KRR) problems with a moderate number of data points ($10^{4}\leq N\leq 10^{7}$). The proposed algorithms based on RPCholesky and KRILL preconditioning are relatively fast, robust, and reliable, outperforming previous approaches [17, 25, 26, 57, 47, 48, 34].
@@
Eigenvalue decay is the main requirement limiting the performance of RPCholesky preconditioning, and this limitation becomes apparent when the regularization $\mu$ is small (see the right panel of Figure˜1). Future work is needed to discover effective preconditioners for KRR problems with slow eigenvalue decay. Methods based on hierarchical decompositions [17, 3, 12, 60] and sparse inverse approximations [50, 28, 61] are promising candidates for this purpose, but there is a concern these methods perform poorly in high dimensions [3].
@@
With appropriate parameter choices and neglecting rounding errors, KRILL is guaranteed to control the condition number regardless of the regularization and the kernel matrix eigenvalue decay. Nonetheless, there remain opportunities to improve on KRILL. First, the existing analysis of sparse random sign embeddings [16] does not specify appropriate values of $\zeta$ and $d$ to use in practice. More analysis of sparse sign embeddings is needed to close the theory–practice gap; recent papers [14, 54] make some progress in this direction. A second potential concern with KRILL is numerical stability; see the recent papers [36, 23]. As a practical remedy, we have addressed numerical stability by adding a small shift to the regularizer, but this approach has not been analyzed rigorously. Finally, it may be possible to develop more efficient algorithms for restricted KRR problems when the number of data centers is $k\gg N^{1/2}$. The ideal conjugate gradient algorithm would require $\mathcal{O}(kN)$ operations, but KRILL is more expensive due to the cost of inverting the $k\times k$ preconditioner.
@@ Acknowledgments
We would like to acknowledge helpful conversations with Misha Belkin, Tyler Chen, Jackie Lok, Riley Murray, Madeleine Udell, and Jorge Garza-Vargas.
@@ Appendix A Data sets
Data set Dimension $d$ Sample size NN Source